Hot Teas common stock is currently selling for $38.97. The last annual dividend paid was $1.26 per share and the market rate of return is 13.2 percent. At what rate is the dividend growing?
Answer #1

ke = Market rate of return = 13.2%

P0 = 38.97

D0 = 1.26

ke = D0*(1+g)/P0 + g

0.132 = 1.26*(1+g)/38.97 + g

0.132 - g = (1.26/38.97)*(1+g)

0.132 - g = 0.03233*(1+g)

0.132 - g = 0.03233 + 0.03233(g)

0.09967 = 1.03233 (g)

g = 0.09967/1.03233 = 0.0965 = 9.65%

Answer : 9.65%